OpenStudy (anonymous):

Convert the rectangular equation to a polar equation. Express r in terms of "theta". 7x-4y=1

OpenStudy (anonymous):

@satellite73 @Loser66 please

zepdrix (zepdrix):

When converting from rectangular to polar, we can use the fact that:\[\large x=r \cos \theta \qquad \qquad \qquad y=r \sin \theta\] Sooooo,\[\large 7x-4y=1 \qquad\to\qquad 7r \cos \theta-4r \sin \theta=1\]Factoring an r out of each term gives us,\[\large r(7\cos \theta-4 \sin \theta)=1\]Then to solve for r we simply divide that big messy thing to the other side. Is this what you're trying to do with this problem? :o

OpenStudy (anonymous):

Yes, I think that is what I need to do. Trig confuses me so much. so: \[r(7\cos \Theta-4\sin \theta)=1\] How do I write that in Polar form?

zepdrix (zepdrix):

Dividing each side by \(\large 7\cos\theta-4\sin\theta\) gives us, \[\large r=\frac{1}{7\cos \theta-4\sin \theta}\] I think this is the polar form they're asking for, with r expressed as a function of theta. Which is what the instructions were asking for: `Express r in terms of "theta".`

zepdrix (zepdrix):

Trig confuses you??? :) Aw man that's my favorite!

OpenStudy (anonymous):

Okay thank you! Yeah, it's probably because I did summer trig so it's a little too fast for me. :(

OpenStudy (anonymous):

I'm studying for my final right now. so far I have a low B, I'm just trying to keep from getting a C. thank you again for your help!

zepdrix (zepdrix):

oh neato c: